OpenStudy (anonymous):

Percy paid $24.10 for a basketball. The price of the basketball was $22.99. What was the sales tax rate?

hero (hero):

Any idea on how to get started with this one?

OpenStudy (anonymous):

no i don't get it yet.

hero (hero):

Scenario: Percy purchased a basketball. The ORIGINAL PRICE of the ball was $22.99. But the Actual AMOUNT PAID (SALES PRICE) was $24.10. Notice that $24.10 is larger than the original price of $22.99. This means the original price was increased by a certain amount known as the sales rate. We need to figure how much sales rate was. To figure out what the sales tax rate is, we use a simple formula: Original Price plus a certain percentage of the original price equals the sales price. Written mathetmatically: x + %x = y x = original price y = sales price % = sales rate In this case: x = 22.99 y = 24.10 % = unknown So inputting all the known values into the formula, we get: 22.99 + %(22.99) = 24.10 Now we solve for %, which is the sales rate: 22.99 + %(22.99) = 24.10 Subtract 22.99 from both sides: %(22.99) = 24.10 - 22.99 %(22.99) = 1.11 Divide both sides by 22.99: % = 1.11/22.99 % = 4.8 So the sales rate is approximately 4.8 percent.

hero (hero):

I just wanted to add more description to the scenario which actually attempts to explain what's going on and exactly what it is we need to find.

OpenStudy (anonymous):

thanks a lot...Hero

hero (hero):

You actually understood it?

OpenStudy (anonymous):

yes

hero (hero):

That what I provided was a mouthful. I'd be surprised if you actually understood all of that on the first go.

OpenStudy (anonymous):

Ethan will be going on a study tour to England in a week. Which currency should he use when he is in England?

OpenStudy (anonymous):

my hero !

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Tanya wants to order 50 pizzas for a party. However, the pizza supplier can deliver only 45 pizzas on the given date. How will you describe the relationship between demand and supply of pizzas? A. Demand is equal to supply. B. Demand is greater than supply. C. Demand is less than supply.

OpenStudy (anonymous):

anyone there can reply to me the answer?